OpenStudy (anonymous):

In a clinical trial is used reduce fever cause side effects and 6% of patients who took it . Three patients are selected at random find the probability that all of the patients had side effects

ganeshie8 (ganeshie8):

probability for selecting one patient wid side effects = 0.06

ganeshie8 (ganeshie8):

so probability for selecting 3 patients wid side effects = 0.06 * 0.06 * 0.06 = (0.06)^3 = ?